The perimeter of a square is twice the perimeter of an equilateral triangular lot. If one side of the square is 60 meters, find the length of a side of the triangle.
Answer by amoresroy(361)

The perimeter of a square is twice the perimeter of an equilateral triangular lot. If one side of the square is 60 meters, find the length of a side of the triangle.
Let S = perimeter of a square
S/4 = side of square
T = perimeter of an equilateral triangular lot
T/3 = length of side of equilateral triangle
S = 2T
If S/4 = 60
S = 240
T = S/2 = 120
T/3 = 120/3 = 40
The length of the side of the triangle is 40 meters